# Intermittent time series forecasting with Gaussian Processes and Tweedie likelihood

### v4

@@ 3.2 The Tweedie likelihood
The Tweedie is a family of exponential dispersion models (a generalization of the exponential family, see Dunn and Smyth, 2005) characterized by a power mean-variance relationship. A non-negative random variable $Y$ is distributed as a Tweedie, $Y\sim\mathrm{Tw}(\mu,\phi,\rho)$, if $Y$ follows an exponential dispersion model distribution such that
@@
Depending on $\rho$, the Tweedie family includes the Gaussian ($\rho=0$), Poisson ($\rho=1$), and Gamma ($\rho=2$) distributions. However, we restrict $\rho\in(1,2)$ to have a non-negative distribution with a possible point mass in zero. In this setting, the Tweedie (Dunn and Smyth, 2005) is a Poisson mixture of Gamma distributions:
@@
where $\mathbb{P}$ denotes the probability of an event and $\mathrm{Ga}(\cdot)$ is the density of a Gamma distribution. The Tweedie distribution is generally bimodal (Fig. 4): the first mode is in 00, while the second one is the mode of the mixture of Gamma distributions. This mixture has a continuous density function on the positive real values. Count forecasts can be generated by rounding the samples.
The parametrization (7-8), though interpretable, is not usable as a likelihood function. Indeed there is no theoretical result for truncating the infinite sum in eq. (8) while controlling the exceedance probability. In order to evaluate the Tweedie we use the $(\mu,\phi,\rho)$ parametrization (Dunn and Smyth, 2005). The probability of zero and the continuous density for $y>0$ are:
@@
where $\alpha$ is defined in eq. (6). Dunn and Smyth (2005) provide a truncating rule for evaluating the infinite summation of eq. (11). It exploits the fact that $V(j)$ is a concave function of $j$. In  B.1 we report the derivations of the truncating rule and we show that we generally require less than 10 terms in the summation of eq. (11) for a precise approximation. The Tweedie density is complex to implement; yet, its evaluation only requires a small overhead compared to other distributions.
@@
where $\boldsymbol{\theta}_{\mathrm{lik}}=\{\phi,\rho\}$. We set $\mu=\text{softplus}(f)$: the latent variable affects, through $\lambda$ and $\beta$, both the mass in 00 and the distribution on the positive $y$, see eq. (6), (7), (8). Thus, a bimodal forecast distribution parametrized by a single latent process. The hyper-parameters $\phi$ and $\rho$ are optimized and they are equal for all the data points in the same time series.
@@ 3.2.1 Comparison with Tweedie loss
Januschowski et al. (2022) mention the use of the Tweedie loss for tree-based models as one of the reasons of success in the M5 competition. Also Jeon and Seong (2022) obtained good results in the M5 competition by training a DeepAR (Salinas et al., 2020) with the Tweedie loss. The Tweedie loss is indeed available both in lightGBM [^1] https://lightgbm.readthedocs.io/en/latest/Parameters.html and in PyTorch Forecasting [^2] https://pytorch-forecasting.readthedocs.io/en/stable/api/pytorch_forecasting.metrics.point.TweedieLoss.html. But such models only return point forecasts. We clarify, for the first time, that the Tweedie loss is obtained from the Tweedie distribution via a rough approximation, which makes it unsuitable for probabilistic forecasting.
The Tweedie loss (Jeon and Seong, 2022) is:
@@
which approximates the Tweedie density in eq. (10) by setting both $A{\color[rgb]{0.,0.,0.}(y)}=1$ and $\phi=1$. Setting $A{\color[rgb]{0.,0.,0.}(y)}=1$ for all $y>0$ avoids evaluating the infinite summation in eq. (11); however, this makes the distribution unimodal and shortens its tails. Indeed, $A$ is a function of $y$ and not a simple normalization constant. Moreover, fixing the dispersion parameter to $\phi=1$ further reduces the flexibility. In Sec. 4.5 we show that the accuracy of TweedieGP worsens when we adopt the approximation with $A{\color[rgb]{0.,0.,0.}(y)}=1$, $\phi=1$ rather than the actual Tweedie distribution.
@@ 4 Experiments
We test the performance of our models on 5 publicly available data sets. We consider a time series as intermittent if the average demand interval (ADI) is strictly greater than 1, that is, all the time series containing at least one observation equal to zero (Syntetos et al., 2005). The data sets are:
In Tab. 1 we report the length of the training data ($T$) and of the forecast horizon ($h$) of each data set; overall we consider more than 40’000 intermittent time series. Fig. 5 shows how the proportion of zeros and the demand size varies across the time series of each data set.
The data sets with the highest proportion of zeros (median $0.9$) are RAF and OnlineRetail (Fig. 5 left). However, the time series of OnlineRetail are about five times longer than those of RAF, Tab. 1. In contrast, Auto has the lowest proportion of zeros and its time series are also the shortest ($T=18$). The M5 data set is the most heterogeneous as for proportion of zeros and it generally has narrow demand size. It also contains the largest amount (30490) of time series, which are also very long ($T=1941$).
@@ 4.1 GP training
We implement our models in GPyTorch (Gardner et al., 2018). We estimate the variational parameters, the hyper-parameters of the GP ($c,{\color[rgb]{0.,0.,0.}\ell},\sigma^{2}$) and of the likelihood (p for the negative binomial distribution, $\phi$ and $\rho$ for the Tweedie) via gradient descent for 100 iterations with early stopping, using Adam optimizer (Kingma and Ba, 2017) with learning rate $0.1$.
@@ 4.2 Baselines
We compare NegBinGP and TweedieGP against empirical quantiles (EmpQuant), WSS, ADIDAC, and iETS. We implemented on our own WSS (Willemain et al., 2004), drawing 50’000 samples from its forecast distribution.
We use ADIDA (Nikolopoulos et al., 2011) generating prediction intervals via conformal inference (Angelopoulos and Bates, 2023), using the implementation provided by statsforecast (Garza et al., 2022). We refer to it as ADIDAC, where the subscript stands for conformal.
We use iETS (Svetunkov and Boylan, 2023) from the smooth package (Svetunkov (2024), v. 4.1.1). We use the most complete model (iETSA) which fits four different occurrence models and selects among them via AICc. We draw 50’000 samples from the forecast distribution.
In D, we show how to reproduce our experiments and we detail the integration of our models in GPyTorch (Gardner et al., 2018).
We do not include global models such as LightGBM (Ke et al., 2017), which performed well on the M5 competition (Januschowski et al., 2022), since we focus our comparison on local models. We leave to future works the comparison between local and global models.
@@ 4.3 Metrics
Denoting by $\hat{y}^{q}$ the forecast of quantile $q$, and by $y$ the observed value, the quantile loss (Gneiting and Raftery, 2007) is:
@@
The quantile loss is however scale-dependent. In order to obtain a scale-free indicator (Athanasopoulos and Kourentzes, 2023) we scale it by the quantile loss of the empirical quantiles ($\mathrm{emp}^{q}$) on the training data, and we average it through the forecast horizon. Let $\hat{\mathbf{y}}^{q}_{T+1:T+h}:=(\hat{y}_{T+1}^{q},\dots,\hat{y}_{T+h}^{q})^{\top}$ be the forecasts of quantile $q$ from times $T+1$ to $T+h$; the scaled quantile loss is:
@@
The scaled quantile loss scores the quantile predictions as point forecasts. We then assess the entire predictive distribution, using an indicator similar to the Ranked Probability Score (RPS, Czado et al., 2009). The RPS can be approximated by averaging the quantile loss computed for multiple probability levels (Bracher et al., 2021). For many applications involving intermittent time series, the lower part of the forecast distribution is not relevant (Boylan and Syntetos, 2006). We thus focus on a set of probability levels $L=\{0.5,0.55,0.6,0.65,0.7,0.75,0.8,0.85,0.9,0.95,0.99\}$ with cardinality $|L|=11$ and we propose the metric
@@
We also report the Root Mean Squared Scaled Error (RMSSE) (Hyndman and Koehler, 2006) of the point forecasts:
@@
where $\hat{\mathbf{y}}_{T+1:T+h}:=(\hat{y}_{T+1},\dots,\hat{y}_{T+h})^{\top}$ denotes a vector of forecasts. We use the mean of the forecast distribution as point forecast, as this is the minimizer of RMSSE (Kolassa, 2020).
@@ 4.4 Discussion
$\mathrm{sQ}_{0.8}$ 1.71 1.74 1.43 1.61 1.45 1.44
$\mathrm{sQ}_{0.9}$ 1.54 1.62 1.33 1.53 1.28 1.25
$\mathrm{sQ}_{0.95}$ 1.38 1.49 1.34 1.58 1.18 1.15
$\mathrm{sQ}_{0.99}$ 1.24 1.39 1.65 2.20 1.16 1.15
@@
$\mathrm{sQ}_{0.8}$ 2.33 2.33 2.28 2.39 2.26 2.25
$\mathrm{sQ}_{0.9}$ 2.31 2.29 2.23 2.50 2.23 2.20
$\mathrm{sQ}_{0.95}$ 2.33 2.29 2.34 2.75 2.29 2.22
$\mathrm{sQ}_{0.99}$ 3.01 2.88 4.08 4.35 3.21 2.97
@@
OnlineRetail rmsse 0.97 0.98 0.94 0.94 0.94 0.95
$\mathrm{sQ}_{0.5}$ 1.19 1.35 1.25 1.20 1.18 1.19
$\mathrm{sQ}_{0.8}$ 1.31 1.47 1.36 1.36 1.30 1.29
$\mathrm{sQ}_{0.9}$ 1.48 1.58 1.58 1.50 1.43 1.42
$\mathrm{sQ}_{0.95}$ 1.80 1.77 2.17 1.71 1.64 1.62
$\mathrm{sQ}_{0.99}$ 4.16 2.55 7.21 2.85 2.47 2.60
@@
$\mathrm{sQ}_{0.8}$ 1.18 1.33 1.14 1.15 1.10 1.09
$\mathrm{sQ}_{0.9}$ 1.25 1.45 1.16 1.24 1.16 1.13
$\mathrm{sQ}_{0.95}$ 1.32 1.63 1.34 1.46 1.17 1.19
@@
We show in Tab. 2 the score of each method, averaged with respect to all the time series of each data set. In each row, we boldface the model with the lowest average loss and the models whose average loss is not significantly different from the best. We test significance of the difference in loss by using the paired $t$-test with FDR correction for multiple comparisons (Hastie et al., 2009, Sec 18.7.1).
We argue that models should be compared considering the mean of the quantile losses, and not a rank-based statistic. The distribution of the quantile loss is strongly asymmetric: most forecasts have a small quantile loss; only a few forecasts incur a large penalty. The mean is sensitive to such rare and large losses, unlike rank statistics. Moreover, the correct predictive distribution is guaranteed to minimize the expected value of the scoring rule (Gneiting and Raftery, 2007), but there are no guarantees for rank statistics.
Scaled quantile loss
@@
iETS is a good match for the GP models on the lowest quantile we consider, but it is generally outperformed on the highest quantiles. Also in Svetunkov and Boylan (2023), iETS is outperformed by other methods on the average quantile loss.
@@
The model ranking remains consistent if we consider as intermittent only the time series with ADI $>1.32$. This cut-off was introduced by Syntetos et al. (2005), studying whether Croston’s method or the Syntetos-Boylan approximation (Syntetos and Boylan, 2005) yields a lower MSE. It is also sometimes used as a cut-off for intermittent demand (Long et al., 2025); the practical effect is that the selected time series are more sparse. In the M5, Carparts, and RAF data sets, this criterion reduced the number of time series by less than 5%, implying only minor changes to the numerical results. The exceptions are the OnlineRetail data set, in which the number of intermittent time series decreases from 2452 to 2023, and Auto, where only 1227 time series out of 3000 are retained. Also in these cases, however, the model rankings do not change, providing favorable evidence for the GP models. The results are reported in Tab. 6.
@@
On the Auto data set, which has moderate intermittency, only TweedieGP has almost correct coverage on all quantiles above $0.5$. A similar comment can be done for M5, even though in this case the proportion of zeros is larger. Correct coverage is provided, only by TweedieGP and empirical quantiles, for quantiles above $0.8$. On the RAF data set, the very sporadic positive demand implies forecast overcoverage apart from the highest quantiles.
@@
M5 $0.10\pm 0.03$ $0.01\pm 0.01$ $0.76\pm 0.07$ $1.20\pm 0.39$ $2.12\pm 0.49$
OnlineRetail $0.11\pm 0.05$ $0.01\pm 0.01$ $0.44\pm 0.06$ $0.32\pm 0.18$ $0.35\pm 0.22$
Auto $0.04\pm 0.02$ $0.01\pm 0.01$ $0.13\pm 0.03$ $0.20\pm 0.06$ $0.20\pm 0.08$
Carparts $0.05\pm 0.02$ $0.01\pm 0.01$ $0.14\pm 0.03$ $0.17\pm 0.11$ $0.24\pm 0.12$
RAF $0.05\pm 0.03$ $0.01\pm 0.01$ $0.20\pm 0.03$ $0.22\pm 0.06$ $0.30\pm 0.07$
@@
We report in Tab. 3 the computational times per time series of each method, including training and generating the forecasts on time series each data set. Empirical quantiles are practically immediate and their execution times are not shown in the table. ADIDAC and WSS are the fastest methods.
The computational time of our GPs is comparable with that of iETS. On the data sets containing the longest time series (M5), the average computational time of the GPs is about 2 seconds. An exact GP implementation would be rather slow on the time series of the M5 data set, due to its cubic computational complexity in the number of observations $T$. Thanks to the variational approximation, the computational complexity of our models is instead quadratic in $T$ for $T<200$. For longer time series the cost is capped by the variational formulation and becomes linear in $T$. See A for more details.

### v5

@@ 3.2 The Tweedie likelihood
The Tweedie is a family of exponential dispersion models (a generalization of the exponential family, see Dunn_Smyth_2005) characterized by a power mean-variance relationship. A non-negative random variable $Y$ is distributed as a Tweedie, $Y\sim\mathrm{Tw}(\mu,\phi,\rho)$, if $Y$ follows an exponential dispersion model distribution such that
@@
Depending on $\rho$, the Tweedie family includes the Gaussian ($\rho=0$), Poisson ($\rho=1$), and Gamma ($\rho=2$) distributions. However, we restrict $\rho\in(1,2)$ to have a non-negative distribution with a possible point mass in zero. In this setting, the Tweedie (Dunn_Smyth_2005) is a Poisson mixture of Gamma distributions:
@@
where $\mathbb{P}$ denotes the probability of an event and $\mathrm{Ga}(\cdot)$ is the density of a Gamma distribution (parametrized by shape and rate). The Tweedie distribution is generally bimodal (Fig. 4): the first mode is in 0, while the second one is the mode of the mixture of Gamma distributions. This mixture has a continuous density function on the positive real values. Count forecasts can be generated by rounding the samples.
The parametrization (7-8), though interpretable, is not usable as a likelihood function. Indeed there is no theoretical result for truncating the infinite sum in eq. (8) while controlling the exceedance probability. In order to evaluate the Tweedie we use the $(\mu,\phi,\rho)$ parametrization (Dunn_Smyth_2005). The probability of zero and the continuous density for $y>0$ are:
@@
where $\alpha$ is defined in eq. (6). Dunn_Smyth_2005 provide a truncating rule for evaluating the infinite summation of eq. (11). It exploits the fact that $V(j)$ is a concave function of $j$. In  B.1 we report the derivations of the truncating rule and we show that we generally require less than 10 terms in the summation of eq. (11) for a precise approximation. The Tweedie density is complex to implement; yet, its evaluation only requires a small additional overhead compared to other distributions.
@@
where $\boldsymbol{\theta}_{\mathrm{lik}}=\{\phi,\rho\}$. We set $\mu=\text{softplus}(f)$: the latent variable affects, through $\lambda$ and $\beta$, both the mass in 0 and the distribution on the positive $y$, see eq. (6), (7), (8). Thus, a bimodal forecast distribution is parametrized by a single latent process. The hyper-parameters $\phi$ and $\rho$ are optimized and they are equal for all the data points in the same time series.
@@ 3.2.1 Comparison with Tweedie loss
januschowski2022forecasting mention the use of the Tweedie loss for tree-based models as one of the reasons of success in the M5 competition. Also jeon2022robust obtained good results in the M5 competition by training a DeepAR (salinas2020deepar) model with the Tweedie loss. The Tweedie loss is indeed available both in lightGBM [^4] https://lightgbm.readthedocs.io/en/latest/Parameters.html and in PyTorch Forecasting [^5] https://pytorch-forecasting.readthedocs.io/en/stable/api/pytorch_forecasting.metrics.point.TweedieLoss.html. But such models only return point forecasts. We clarify, for the first time, that the Tweedie loss is obtained from the Tweedie distribution via a rough approximation, which makes it unsuitable for probabilistic forecasting.
The Tweedie loss (jeon2022robust) is:
@@
which approximates the Tweedie density in eq. (10) by setting both $A{\color[rgb]{0.,0.,0.}(y)}=1$ and $\phi=1$. Setting $A{\color[rgb]{0.,0.,0.}(y)}=1$ for all $y>0$ avoids evaluating the infinite summation in eq. (11); however, this makes the distribution unimodal and shortens its tails. Indeed, $A$ is a function of $y$ and not a simple normalization constant. Moreover, fixing the dispersion parameter to $\phi=1$ further reduces the flexibility. In Sec. 4.5 we show that the accuracy of TweedieGP worsens when we adopt the approximation with $A{\color[rgb]{0.,0.,0.}(y)}=1$ and $\phi=1$, rather than the actual Tweedie distribution.
@@ 4 Experiments
We test the performance of our models on 5 publicly available data sets. We consider a time series as intermittent if the average demand interval (ADI) is strictly greater than 1, that is, all the time series containing at least one observation equal to zero (Syntetos_Boylan_Croston_2005). The data sets are:
In Tab. 1 we report the length of the training data ($T$) and of the forecast horizon ($h$) of each data set; e consider overall more than 40’000 intermittent time series. Fig. 5 shows how the proportion of zeros and the demand size varies across the time series of each data set.
The data sets with the highest proportion of zeros are RAF and OnlineRetail (Fig. 5 left). However, the time series of OnlineRetail are about five times longer than those of RAF, Tab. 1. In contrast, Auto has the lowest proportion of zeros and its time series are also the shortest ($T=18$). The M5 data set is the most heterogeneous as for proportion of zeros and it generally has narrow demand size. It also contains the largest amount (30490) of time series, which are also very long ($T=1941$).
@@ 4.1 GP training
We implement our models in GPyTorch (Gardner_Pleiss_Weinberger_Bindel_Wilson_2018). We estimate the variational parameters, the hyper-parameters of the GP ($c,{\color[rgb]{0.,0.,0.}\ell},\sigma^{2}$) and of the likelihood (p for the negative binomial distribution, $\phi$ and $\rho$ for the Tweedie) via gradient descent for 100 iterations with early stopping, using Adam optimizer (Kingma_Ba_2017) with learning rate $0.1$.
@@ 4.2 Baselines
We compare NegBinGP and TweedieGP against empirical quantiles (EmpQuant), WSS, ADIDAC, and iETS. We implemented on our own WSS (Willemain_Smart_Schwarz_2004), drawing 50’000 samples from its forecast distribution.
We use ADIDA (Nikolopoulos_Syntetos_Boylan_Petropoulos_Assimakopoulos_2011) generating prediction intervals via conformal inference (Angelopoulos_Bates_2022), using the implementation provided by statsforecast (Garza_Canseco_Challú_Olivares_2022, v. 1.7.6). We refer to it as ADIDAC, where the subscript stands for conformal.
We use iETS (Svetunkov_Boylan_2023) from the smooth R package (Svetunkov_smooth, v. 4.1.1). We use the most complete model (iETSA) which fits four different occurrence models and selects among them via AICc. We draw 50’000 samples from the forecast distribution.
In D, we show how to reproduce our experiments and we detail the integration of our models in GPyTorch (Gardner_Pleiss_Weinberger_Bindel_Wilson_2018).
We do not include global models such as LightGBM (lighGBM), which performed well on the M5 competition (januschowski2022forecasting), since we focus our comparison on local models. We leave to future works the comparison between local and global models.
@@ 4.3 Metrics
Denoting by $\hat{y}^{q}$ the forecast of quantile $q$, and by $y$ the observed value, the quantile loss (Gneiting_Raftery_2007) is:
@@
The quantile loss is however scale-dependent. In order to obtain a scale-free indicator (Athanasopoulos_Kourentzes_2023) we scale it by the quantile loss of the empirical quantiles ($\mathrm{emp}^{q}$) on the training data, and we average it through the forecast horizon. Let $\hat{\mathbf{y}}^{q}_{T+1:T+h}:=(\hat{y}_{T+1}^{q},\dots,\hat{y}_{T+h}^{q})^{\top}$ be the forecasts of quantile $q$ from times $T+1$ to $T+h$; the scaled quantile loss is:
@@
The scaled quantile loss scores the quantile predictions as point forecasts. We then assess the entire predictive distribution, using an indicator similar to the Ranked Probability Score (RPS, Czado_Gneiting_Held_2009). The RPS can be approximated by averaging the quantile loss computed for multiple probability levels (Bracher_Ray_Gneiting_Reich_2021). For many applications involving intermittent time series, the lower part of the forecast distribution is not relevant (boylan_syntetos_2006). We thus focus on a set of probability levels $L=\{0.5,0.55,0.6,0.65,0.7,0.75,0.8,0.85,0.9,0.95,0.99\}$ with cardinality $|L|=11$ and we propose the metric
@@
We also report the Root Mean Squared Scaled Error (RMSSE) (Hyndman_Koehler_2006) of the point forecasts:
@@
where $\hat{\mathbf{y}}_{T+1:T+h}:=(\hat{y}_{T+1},\dots,\hat{y}_{T+h})^{\top}$ denotes a vector of forecasts. We use the mean of the forecast distribution as point forecast, as this is the minimizer of RMSSE (kolassa2020best).
@@ 4.4 Discussion
$\mathrm{sQ}_{0.8}$ 1.71 1.74 1.43 1.61 1.46 1.45
$\mathrm{sQ}_{0.9}$ 1.54 1.62 1.33 1.53 1.28 1.26
$\mathrm{sQ}_{0.95}$ 1.38 1.49 1.34 1.58 1.18 1.16
$\mathrm{sQ}_{0.99}$ 1.24 1.39 1.65 2.20 1.16 1.16
@@
$\mathrm{sQ}_{0.8}$ 2.33 2.33 2.28 2.39 2.24 2.25
$\mathrm{sQ}_{0.9}$ 2.31 2.29 2.23 2.50 2.19 2.19
$\mathrm{sQ}_{0.95}$ 2.33 2.29 2.34 2.75 2.24 2.21
$\mathrm{sQ}_{0.99}$ 3.01 2.88 4.08 4.35 3.09 2.98
@@
OnlineRetail
rmsse 0.97 0.98 0.94 0.94 0.94 0.95
$\mathrm{sQ}_{0.5}$ 1.19 1.35 1.25 1.20 1.17 1.19
$\mathrm{sQ}_{0.8}$ 1.31 1.47 1.36 1.36 1.30 1.30
$\mathrm{sQ}_{0.9}$ 1.48 1.58 1.58 1.50 1.42 1.42
$\mathrm{sQ}_{0.95}$ 1.80 1.77 2.17 1.71 1.62 1.63
$\mathrm{sQ}_{0.99}$ 4.16 2.55 7.21 2.85 2.43 2.63
@@
$\mathrm{sQ}_{0.8}$ 1.18 1.33 1.14 1.15 1.09 1.09
$\mathrm{sQ}_{0.9}$ 1.25 1.45 1.16 1.24 1.15 1.13
$\mathrm{sQ}_{0.95}$ 1.32 1.63 1.34 1.46 1.23 1.20
@@
We show in Tab. 2 the score of each method, averaged with respect to all the time series of each data set. In each row, we boldface the model with the lowest average loss and the models whose average loss is not significantly different from the best. We test significance of the difference in loss by using the paired $t$-test with FDR correction for multiple comparisons (hastie2009elements, Sec 18.7.1).
We argue that models should be compared considering the mean of the quantile losses, and not a rank-based statistic. The distribution of the quantile loss is strongly asymmetric: most forecasts have a small quantile loss; only a few forecasts incur a large penalty. The mean is sensitive to such rare and large losses, unlike rank statistics. Moreover, the correct predictive distribution is guaranteed to minimize the expected value of the scoring rule (Gneiting_Raftery_2007), but there are no guarantees for rank statistics.
Scores
@@
iETS is a good match for the GP models on the lowest quantile we consider, but it is generally outperformed on the highest quantiles. Also in Svetunkov_Boylan_2023, iETS is outperformed by other methods on the average quantile loss.
@@
The model ranking remains consistent if we consider as intermittent only the time series with ADI >1.32>1.32. This cut-off was introduced by Syntetos_Boylan_Croston_2005, studying whether Croston’s method or the Syntetos-Boylan approximation (Syntetos_Boylan_2005) yields a lower MSE. It is also sometimes used as a cut-off for intermittent demand (Long_Bui_Oktavian_Schmidt_Bergmeir_Godahewa_Lee_Zhao_Condylis_2025); the practical effect is that the selected time series are more sparse. In the M5, OnlineRetail, Carparts, and RAF data sets, this criterion reduced the number of time series by less than 5%, implying only minor changes to the numerical results. The only exception is Auto, where only 1227 time series out of 3000 are retained. Also in this case, however, the model rankings do not change, providing favorable evidence for the GP models. The results are reported in C.
@@
On the Auto data set, which has moderate intermittency, TweedieGP has almost correct coverage on all quantiles above $0.5$. A similar comment can be done for M5, even though in this case the proportion of zeros is larger. Correct coverage is provided, only by TweedieGP and empirical quantiles, for quantiles above $0.8$. On the RAF data set, the very sporadic positive demand implies forecast overcoverage apart from the highest quantiles.
@@
M5 $\pm 0.03 0.01±0.01$$\pm 0.01$ $0.76\pm 0.07$ $1.20\pm 0.39$ $2.12\pm 0.49$
OnlineRetail $11\pm 0.05 0.0$$01\pm 0.01$ $0.44\pm 0.06$ $0.32\pm 0.18$ $0.35\pm 0.22$
Auto $04\pm 0.02 0.0$$01\pm 0.01$ $0.13\pm 0.03$ $0.20\pm 0.06$ $0.20\pm 0.08$
Carparts $05\pm 0.02 0.0$$01\pm 0.01$ $0.14\pm 0.03$ $0.17\pm 0.11$ $0.24\pm 0.12$
RAF $05\pm 0.03 0.0$$01\pm 0.01$ $0\pm 0.03 0.22±0$$2\pm 0.06$ $0.30\pm 0.07$
@@
We report in Tab. 3 the computational times per time series of each method, including training and generating the forecasts on time series of each data set. Empirical quantiles are practically immediate and their execution times are not shown in the table. ADIDAC and WSS are the fastest methods.
The computational time of our GPs is comparable with that of iETS. On the data sets containing the longest time series (M5), the average computational time of the GPs is about 2 seconds. An exact GP implementation would be rather slow on the time series of the M5 data set, due to its cubic computational complexity in the number of observations $T$. Thanks to the variational approximation, the computational complexity of our models is instead quadratic in $T$ for $T\leq 200$. For longer time series the cost is capped by the variational formulation and becomes linear in $T$. See A for more details.